28 November 2007

RHIC Startup:

Two good stores overnight.  Beam intensities were 0.9E11/0.9E9 for both fills.  This is the first shift with reproducible good fills for physics.  STAR turned off ZDC HV in the middle of second fill.  Rebucketing looks a little worse in the second store.  The second store was lost when a 90% BLAM (Beam Loss Monitor) alarm occurred.  Maintenance day started a little early because of this.  PHENIX notes that backgrounds were high at the beginning of the first store. Would like to figure out how to do a vernier scan with the collimators in, or otherwise study the source of the backgrounds.  

Wolfram and Peggy finished the longitudinal realignment work at injection.

MikeB exercised the stochastic cooling system.  Yellow came on well, but not Blue.  The opening and closing is being tested during today's maintenence day.  Only Yellow is required to work for dAu running since stochastic cooling does functions only for Au.  They would like some time at realistic intensities to tune it up.

Optics measurements were taken.

Work on Yellow feedback ramp.  There were several debunches at injection. 

If there is a large tune shift along the bunch train, then scrubbing would be needed, but currently it looks OK.

Plan for Thursday is to complete the measurements of tune shift along the bunch.  Make a decision about scrubbing.  Stochastic cooling setup in Yellow.  Orbit correction along ramp. 


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020